Solve the system of equations.

x-y+z=-8
x+y+z=-6
x+y-z=4

Answer by vasumathi(46)

Solve the system of equations.
x-y+z=-8
x+y+z=-6
x+y-z= 4
solution:add the first and the third equations
x-y+z = -8
x+y-z = 4
2x= -4 since y and -y
and z and -z cancels
we got 2x = -4
divide this by 2 on both sides
x = -4/2 = -2
add second and third equations
x+y+z = -6
x+y-z= 4
2x + 2y = -2
divide both sides of this by 2
we get
x + y = -1
plug in x = -2 we got above in this
-2 + y = -1
add 2 on both sides
-2+2+y =-1+2
y = 1
now plug in the values of x = -2 and y = 1 in the first equation
x - y + z=-8
-2-1+z = -8
-3 +z = -8
add 3 on both sides
-3 + 3 + z = -8+3
z = -5
so the final answer is
x = -2 : y = 1 : z=-5